OpenStudy (anonymous):

Find the coordinates of the point in the first quadrant at which the tangent line to the curve 7x^3 -xy + y^3 = 0 is parallel to the x-axis.

OpenStudy (anonymous):

Am I correct in that I take the derivative first? If so, what's next? Otherwise, what the heck is going on?

OpenStudy (anonymous):

for the line to be parallel to x-axis means slope = 0. First cuadrant means x>0 and y>0. So find the derivative, set it to 0, and find the values (>0) for x and y

OpenStudy (anonymous):

For the derivative, use implicit derivation: \(y'=\large-\frac{F_x}{F_y}=-\frac{21x^2-y}{3y^2-x}\) Now set it equal to 0: \(\large -\frac{21x^2-y}{3y^2-x}=0\) which means: \(y-21x^2=0\) or \(y=21x^2\) remmember that values \(3y^2-x=0\) are not alowed

OpenStudy (anonymous):

Excellent, thank you very much